OpenStudy (rewardent):

8x-5>1 This is my first time on here. I am trying to take a course online and have gotten behind. need to see this broken down how it would be solved. and checked. detailed explanation so I can learn it would be awesome

OpenStudy (anonymous):

metal please

OpenStudy (anonymous):

> 3/4

OpenStudy (triciaal):

8x-5>1 separate the variable ; maintain the inequality what you do to one side also do to the other and +5 to each side 8x -5 + 5 > 1 + 5 8x >6 divide by the coefficient of the variable to get 1*x for this problem you have 8x so divide by 8 any change to one side do same to the other to keep the balance 8x/8>6/8 x >6/8 6/8 can be simplified by dividing by 2 6/2 = 3 and 8/2 = 4 so 6/8 = 3/4 x > 6/8 same as x > 3/4 to check, pick a value in the solution and do one side of the original problem and see how it relates to the other x > 3/4 let's pick 3 8x-5 = 8(3)-5 = 24-5 =19 8x-5 > 1 when x = 3 19 > 1 so it's in the solution set